Number of turning point of y=(x+4)(x+2)(x-1)(x-3)


1
Expert's answer
2021-11-29T12:04:46-0500
"y=(x+4)(x+2)(x-1)(x-3)"

"y=(x^2+6x+8)(x^2-4x+3)"

"y'=(2x+6)(x^2-4x+3)+(2x-4)(x^2+6x+8)"

"=2x^3-8x^2+6x+6x^2-24x+18"

"+2x^3+12x^2+16x-4x^2-24x-32"

"=4x^3+6x^2-26x-14"

Find the critical number(s)


"y'=0=>4x^3+6x^2-26x-14=0"

"2x^3+3x^2-13x-7=0"

"(2x^3+x^2)+(2x^2+x)-(14x+7)=0"

"x^2(2x+1)+x(2x+1)-7(2x+1)=0"

"(2x+1)(x^2+x-7)=0"

"2x+1=0=>x_1=-\dfrac{1}{2}"

"x^2+x-7=0"

"D=(1)^2-4(1)(-7)=29>0"


"x_2=\dfrac{-1-\sqrt{29}}{2}, x_3=\dfrac{-1+\sqrt{29}}{2}"

Critical numbers:


"-\dfrac{1}{2}, \dfrac{-1-\sqrt{29}}{2},\dfrac{-1+\sqrt{29}}{2}"

If "x<\dfrac{-1-\sqrt{29}}{2}, y'<0, y" decreases.


If "\dfrac{-1-\sqrt{29}}{2}<x<-\dfrac{1}{2}, y'>0, y" increases.


If "-\dfrac{1}{2}<x<\dfrac{-1+\sqrt{29}}{2}, y'<0, y" decreases.


If "x>\dfrac{-1+\sqrt{29}}{2}, y'>0, y" increases.


Then the function "y=(x+4)(x+2)(x-1)(x-3)" has "4-1=3" turning points.